GreenHab Report – February 2th

GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 77.5 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 12.68 gallons

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank: 142.33 gallons

Time(s) of watering for crops: 11:56 am

Changes to crops: N/A

Narrative: Continued to conduct daily watering and monitoring of plants and crops in the GreenHab. The second list of plants to potentially harvest was also completed and will be conducted in the coming two days along with another annual interior GreenHab clean-up.

Harvest: N/A

Support/supplies needed: N/A

Journalist Report – February 2th

Crew 274 Journalist Report 2-2-2023

Sol: 4

Author: Tony DiBernardo, Crew Journalist

Title: Steal My Sunshine

Crew 274 woke up this morning at 7:30am to the song “Steal My Sunshine” by LEN and a recorded message from the Grad Studies Office, offering well wishes on their mission. The team made oatmeal and continued their Emotional Recognition Test and Psychological Surveys.

EVA Activity

Today didn’t go entirely as planned for the crew’s scheduled EVAs. Every night, the crew writes their plans and submits them to Mission Control for approval, which was received before the end of our comms window. The crew woke up and began preparing for the EVA, but due to a miscommunication between Mission Control and the crew, the EVA was canceled, giving the crew more time in the hab than originally planned. They planned the upcoming EVAs, debated the possibility of life in the universe, and gave the hab a good cleaning.


After an active first sol, planting all of his micro greens and LDEF seeds, most of GreenHab Officer Hines’ daily activities include recording data and notes on the continual progress of both subjects. Overall, the most significant progress was seen in the rapid development of the cress, with strong evidence of early germination of the broccoli, red cabbage, and arugula micro greens. Unfortunately, no progress has been evident in the LDEF seeds, though they will continue to be monitored closely in the coming days following an adjustment in growth tent temperature and reduction in daily watering to avoid over saturation.


Astronomers Loy and Pena observed multiple features across the Sun’s upper atmosphere, the chromosphere. They observed 8 solar dark spots and multiple granule features across the chromosphere, but the astronomers’ search for prominences came up short with no findings. Photo and video data was not collected today as remaining images from Sol day 3 are still being stacked, analyzed, and configured.

Hab Design

Hab Specialist Bill O’Hara continued the development of a case study of the Mars Desert Research Station. This study will review the architecture of the MDRS habitable elements from the point-of-view of a crew member and discusses the design, activity allocations, use-of-space, traffic flow, outfitting, systems and stowage capability of each. Today O’Hara completed a review of the Science Dome module, Musk Observatory module, and the tunnel system that connects all the modules together. His next steps will include a review of the lower and upper levels of the hab module which will be scheduled around EVA support and participation and other science and operations support.

Resource Utilization in Spaceflight

Commander Guthrie is on sol 4 of a two-week case study on the self-sufficiency of food requirements for short-term space flight. Its goal is to develop a low impact and sustainable dietary requirement for astronauts who may travel as long as two years in space. It hopes to provide an optimal mass to nutrient ratio based on the astronaut. It uses only pre-packaged dehydrated meals curated for the specific nutritional needs of the person. This study was performed at other analogs but this is the first time a woman is serving as the case study subject. As women continue to make a larger presence in spaceflight, it is important to strive to continue to understand food impact requirements. A notable difference from the analog environment and spaceflight, is that many analogs focus on community cooking and meal times whereas analyzing self-sustainment is more analogous to the spaceflight environment.

Tonight, the crew is keeping an eye out for cloudy skies that could affect viewing conditions for astronomical observations, as well as looking forward to a new day to conduct more research and to improve our analog operations as a crew.

Sol Summary Report – February 2th

Crew 274 Sol Summary Report 02-02-2023

Sol: 4

Summary Title: Holding Pattern

Author’s name: Sarah E. Guthrie (“Ceres”), Commander

Mission Status: Nominal

Sol Activity Summary: The day was spent in a holding pattern. Before the crew was able to egress the habitat for EVA today, it was terminated by Mission Control. Therefore the crew spent the day rewriting their EVA proposal for the logistics cache research project, which is an intense study that requires meticulous efforts to plan and prepare for in order to execute. The MDRS Robotic Telescope is still malfunctioning and (again) did not take any photos despite following the troubleshooting guidance provided. The Greenhouse is lush and full of life, it’s a peaceful oasis for those who want to take in sunshine and smell the greenery. As a large crew of eight members (seven men, two women) living in close quarters has been an enjoyable experience and despite being unable to complete several science projects morale is still HIGH, and the crew is full of laughs.

Look Ahead Plan: The crew looks forward to possible cache EVA deployment and new solar images from the astronomy crew (fingers crossed).

Anomalies in work: None

Weather: Sunny, -4C High, -7C Low

Crew Physical Status: Average

EVA: None

Reports to be filed: Sol Summary, Operations, EVA Request, HSO, GHO Report, and Astronomy Report.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site